Suppose 𝐴 and 𝑃 are 3 Γ— 3 matrices and 𝑃-1 exists. If 𝑃-1𝐴𝑃 = ( 1 2 3

0 4 5

0 0 6 ) then find all eigen values of 𝐴2 .


1
Expert's answer
2021-10-19T16:57:01-0400

Let "=A=\begin{bmatrix}
 1&2& 3 \\
 0 & 4 & 5\\0 &0&6
\end{bmatrix}"



A "=\begin{bmatrix}
 1 & 2 & 3 \\
 0 & 4 & 5\\0 & 0 & 6
\end{bmatrix}" "=1\begin{vmatrix}
 4 & 5 \\
 0 & 6
\end{vmatrix}-2\begin{vmatrix}
 0&5\\
 0 & 6
\end{vmatrix}+3\begin{vmatrix}
 0 & 4 \\
 0 & 0
\end{vmatrix}"


"=1(24-0)-2(0-0)+3(0-0)"

A"=24\mathrlap{\,/}{=}0"